OpenStudy (anonymous):

Bridget has a kitchen canister, in the shape of a cylinder, which holds flour. If the canister is 22 centimeters wide and 24 centimeters tall, what is its approximate surface area in terms of pi? 770π square centimeters 816π square centimeters 528π square centimeters 1,012π square centimeters

OpenStudy (geometry_hater):

lol i looked up the answer i got it but no explanation

OpenStudy (geometry_hater):

Do you know the formula for getting surface area?

jimthompson5910 (jim_thompson5910):

Hint: Surface area of a cylinder (SA) SA = 2*pi*r^2 + 2*pi*r*h where 'r' is the radius and 'h' is the height

jimthompson5910 (jim_thompson5910):

In your case, r = 22/2 = 11 and h = 24

OpenStudy (anonymous):

thnks so much

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

my answer is b